OpenStudy (anonymous):

the domain of 2/x+1 -3

OpenStudy (anonymous):

is that (2/x)+1-3, or is that (2/x+1)-3?

OpenStudy (anonymous):

(2/x+1) -3

OpenStudy (anonymous):

k. so set it equal to 0. You get (2/x+1)-3=0, move three to the other side, and get (2/x+1)=3. Then cross multiply and you get 2=3(x+1). Foil it out. You get 2=3x+3. Solves for -1/3=x. Does that make sense?

OpenStudy (anonymous):

yeah got it so the domain is -1/3

OpenStudy (anonymous):

ooohhh. No its not. my bad ok so

OpenStudy (anonymous):

(2/x+1)-3. Ok, what number can you NOT divide by in math?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

You CANT divide by 0 because it doesnt work. So set the denominator in 2/(x+1), which is x+1, equal to 0. So you get x+1=0. Move 1 to the other side and you get x=-1. So the domain of the function is from

OpenStudy (anonymous):

its -1 ?

OpenStudy (anonymous):

-infinity to -1, -1 to +infinity

OpenStudy (anonymous):

Yes its -1

OpenStudy (anonymous):

because the denominator in your question is x+1. 3 is not a part of it

OpenStudy (anonymous):

and if you want that domain in range notation, its -infinity<-1<+infinity i believe

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

so the domain for (-3/x+4 )+2 is -2 ?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

if you have x+4 as your denominator, set that equal to zero. what do you get?

OpenStudy (anonymous):

and no its not. The domain can be everything BUT what makes the domain 0. So if it cant be -1, domain is everything else from -infinity to +infinity

OpenStudy (anonymous):

If you want, I can do a more in depth answer so I explain everything. Would you prefer that?

OpenStudy (anonymous):

its actually x-4

OpenStudy (anonymous):

Ok. So what is domain? Domain is values of x. Its easier to find everything that x CANT be. In the case of the equation you have, you are dividing by x. In math, you cant divide by 0, so you set x-4=0. Move that over, you have x=4, so the domain is everything BUT that, from -inifnity to 4, and from 4 to +infinity.

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

Kk! Glad I could help! Sorry if I was a bit muddled at the beginning.

OpenStudy (anonymous):

now the range (2-x-3 )-5

OpenStudy (anonymous):

okay. well I assume that (-2-x-3)-5 is equal to y, so you have (-2-x-3)-5=y, so you can have y be anything from negative to positive infinity. can you see why?

OpenStudy (anonymous):

2/x-3) -5 actually sorry

OpenStudy (anonymous):

kk. what do YOU think it is?

OpenStudy (anonymous):

-infinity 3 to + infinity 3

OpenStudy (anonymous):

remove that last 3 and you have it correct

OpenStudy (anonymous):

(1/x+1 ) -4 +infinity 1 to - infinity

OpenStudy (anonymous):

i think?

OpenStudy (anonymous):

-1 :)

OpenStudy (anonymous):

-1 where?

OpenStudy (anonymous):

because x+1=0, x=-1 if you rearrange

OpenStudy (anonymous):

-1 in your domain

OpenStudy (anonymous):

+infinity -1 to - infinity ?

OpenStudy (anonymous):

yup